media services charges $30 for a phone and $20/month for its economy plan.the equation c=20t+30 describes the total cost,c, of operating a meia service phone for t months.
The total cost of service is ?
If a customer has only $80 available how many months of service can she recieve?
A.3
B.2 1/3
C.2
D. 2 1/2

Answer by nerdybill(7384)

Given:
c=20t+30
c = 80
.
80=20t+30
50=20t
50/20 = t
5/2 = t
2 1/2 = t
.
Answer: D